OpenStudy (anonymous):

1. How is the density of a substance calculated? 2. If the shape of an object is irregular, like a stone, how can you measure its volume? 3. If the volume of a cube of wood is15.625 cm3 and its mass is 10.00 grams, show the calculation for the density of the wood. 4. Based on the density of the wood that you calculated above, would the wood float or sink if it is placed in a bowl of water? Explain your answer. Remember that the density of liquid water is 1.00 g/cm3. 5. A marble or small sphere, made of stone, is found to have a mass of 3.05 grams.

OpenStudy (anonymous):

To find its volume, it is placed in a graduated cylinder that originally contains 10.00 mL of water. When the marble is placed in the graduated cylinder, the new volume of the water and marble

OpenStudy (anonymous):

^ is the rest of part 5

OpenStudy (anonymous):

@hartnn

OpenStudy (anx):

1). the formula is p = m/v where p = density m= mass v = volume. 2). to find the density of an irregular object, find its mass then fill a gradulated cylinder and see what the displacement is. So if you fill 10mL and when you add the rock, its at 14mL 14-10 = 4. 3). p =m/v m = 10.00 v = 15.625 10.00/15.625 = .64 (double check this, did it in my head) 4). The object will float, read up on buoyancy for a futher explanation. http://scienceprimer.com/buoyancy 5). I don't understand the question? What are you trying to find.